A radar signal took 5.1s to go to mars and return. How far away was Mars at that moment


1
Expert's answer
2021-01-26T13:15:46-0500

The distance Earth-Marth is equal

"d=\frac{ct}{2}=\frac{3\times 10^8\rm m/s\times 5.1\: s}{2}=7.65\times 10^8\; \rm m"